a cheetah traveling at 7.0 m/s East speeds up to 22 m/s East in 1.7 s. what is its acceleration?


1
Expert's answer
2021-02-24T12:48:58-0500
"a=\frac{v_f-v_i}{t}\\a=\frac{22-7}{1.7}=8.8\frac{m}{s^2}"